OpenStudy (mony01):

A certain radioactive substance has a half-life of 5 hours. Find the time for a given amount to decay to one-tenth of its original mass

OpenStudy (anonymous):

We want to find the half life so the base of the exponential equation should be 0.5 for simplicity. We also have to divide by the half life (t), and x represents the time. \[y=0.5^{x/t}\] Do you think you know what to do now that you have this equation?

OpenStudy (mony01):

not really

OpenStudy (mony01):

@Riddellikins so x is 5 and what is t?

OpenStudy (wolf1728):

We'll call beginning amount 100% and ending amount 10% elapsed time = half-life * log (beg amt/end amt) / log 2 elapsed time = 5 hours * log (100/10)/0.3010299957 elapsed time = 5 * 1/0.3010299957 elapsed time = 5 * 3.3219280945 elapsed time = 16.6096404724 hours formulas and calculator here: http://1728.org/halflife.htm

OpenStudy (wolf1728):

mony01 you get the same answer?